If the marginal propensity to consume is 2/3, then the government purchases multiplier is
Use letters in alphabetical order to select options
A
2
B
0.33
C
1.67
D
3
Answer #1

Answer D 3
Explanation
Multiplier = 1/MPS
MPS = 1 - MPC = 1 - 2/3 = 1/3

Multiplier = 1/(1/3) = 3